OpenStudy (anonymous):

to how much water should 50. mL of 12 M hydrochloric acid be added to produce a 4.0 M solution?

OpenStudy (anonymous):

\[m _{1} \times v _{1} = m _{2} \times v _{2}\]where m=molarity and v=volume So you have 12M for m1 and 50 mL for v1 and you have 4.0M for m2 and you want to solve for v2. So you can move the equation around and you will get \[\frac{ m _{1} \times v _{1} }{ m _{2} }= v _{2}\] Make sure that you pay attention to your units ... molarity is moles per LITER so your going to need to convert 50 mL to L.

OpenStudy (anonymous):

\[\frac{ .05L \times 12.0M }{ 4.0M }=.15L\]